# Quentin Le Houérou    Ludovic Levy Patey    Keita Yokoyama

### v1

@@ Abstract
In this article, we prove that Ramsey’s theorem for pairs and two colors is a $\forall\Pi^{0}_{4}$ conservative extension of $\mathsf{RCA}_{0}+\mathsf{B}\Sigma^{0}_{2}$, where a $\forall\Pi^{0}_{4}$ formula consists of a universal quantifier over sets followed by a $\Pi^{0}_{4}$ formula. The proof is an improvement of a result by Patey and Yokoyama [21] and a step towards the resolution of the longstanding question of the first-order part of Ramsey’s theorem for pairs.
@@ 1 Introduction
The proof of the main theorem follows the structure of Patey and Yokoyama [21], using an indicator for $\forall\Pi^{0}_{4}$ conservation defined by Yokoyama [26, Section 4].
@@ 1.3 Structure of this paper
Then, in Section 3, we develop a framework for $\forall\Pi^{0}_{4}$ conservation of Ramsey-like-$\Pi^{1}_{2}$ theorems. For this, given a Ramsey-like-$\Pi^{1}_{2}$ theorem $\Gamma$, we first define in Section 3.1 a notion of $n$-density$(\Gamma)$, which informally asserts the existence of sufficiently large finite sets, such that $n$ consecutive applications of $\Gamma$ yield a large set. Following techniques initially defined by Kirby and Paris [14], we prove that $\mathsf{RT}^{2}_{2}$ is a $\forall\Pi^{0}_{4}$ conservative extension of $\mathsf{RCA}_{0}+\mathsf{B}\Sigma^{0}_{2}+$ “there exists $n$-dense sets for every $n\in\omega$” (see Theorem 3.7). $\forall\Pi^{0}_{4}$ conservation over $\mathsf{RCA}_{0}+\mathsf{B}\Sigma^{0}_{2}$ is then reduced to whether $\mathsf{RCA}_{0}+\mathsf{B}\Sigma^{0}_{2}$ proves the existence of these $n$-dense$(\Gamma)$ sets, for every $n\in\omega$. It is however difficult to directly prove the existence of $n$-dense$(\Gamma)$ sets, and we therefore resort in Section 3.2 to the quantitative notion of largeness, $\omega^{n}$-largeness$(T)$, to prove this existence step by step, by handling one application of $\Gamma$ at a time.
@@ 3 A framework for $\forall\Pi^{0}_{4}$ conservation
We now develop a framework for proving that Ramsey-like statements are $\forall\Pi^{0}_{4}$ conservative over $\mathsf{RCA}_{0}+\mathsf{B}\Sigma^{0}_{2}$. This will be divided into two sections : First, building up on the work of Kirby and Paris [14], we prove in Section 3.1 that every Ramsey-like-$\Pi^{1}_{2}$ theorem $\Gamma$ is $\forall\Pi^{0}_{4}$ conservative over $\mathsf{RCA}_{0}+\mathsf{B}\Sigma^{0}_{2}+$ a first-order theory stating the existence of dense$(\Gamma)$ sets. Then, in Section 3.2, we use the quantitative notion of largeness defined in Section 2 to prove the existence of dense$(\Gamma)$ sets, assuming the existence of a combinatorially simpler objects.
The following notion of Ramsey-like-$\Pi^{1}_{2}$ theorem was defined by Patey and Yokoyama [21].
Definition 3.1 (Ramsey-like-$\Pi_{2}^{1}$-formula).
A Ramsey-like-$\Pi_{2}^{1}$-formula is a $\Pi_{2}^{1}$-formula of the form
@@
where $n,k\in\omega$ and $\Psi(f,Y)$ is of the form $(\forall G\subseteq_{fin}Y)\Psi_{0}(f{\upharpoonright}[G]^{n},G)$ with $\Psi_{0}$ a $\Delta_{0}^{0}$-formula.
3.1 Density$(T)$ for Ramsey-like statements
There exists a very general way to characterize the first-order part of a second-order theory, using the notion of density, originally defined by Kirby and Paris [14]. This was later adapted by Bovykin and Weiermann [1, Theorem 1] to prove $\Pi^{0}_{2}$ conservation of combinatorial theorems, and generalized by Patey and Yokoyama [21, Theorem 3.4] for $\forall\Pi^{0}_{3}$ conservation.
@@ 3.1 Density$(T)$ for Ramsey-like statements
Fix a Ramsey-like-$\Pi_{2}^{1}$-statement
@@
Let $\Gamma$ be a Ramsey-like-$\Pi_{2}^{1}$-statement. Given a countable nonstandard model $\mathcal{M}\models\mathsf{I}\Sigma^{0}_{1}$ and an $M$-finite set $Z\subseteq M$ which is $c\mbox{-dense}(T,\Gamma)$ for some $c\in M\setminus\omega$, then there exists an initial segment $I$ of $M$ such that $(I,\mathsf{Cod}(M/I),a^{\mathcal{M}},A^{\mathcal{M}}\cap I)\models\mathsf{WKL}_{0}+\mathsf{B}\Sigma^{0}_{2}+\Gamma+T$ and $I\cap Z$ is infinite in $I$.
@@
Let $\Gamma$ be a Ramsey-like-$\Pi_{2}^{1}$-statement of the form
@@
Case $n>0$. Suppose the property to be true at rank $n-1$ and fix $Y$ an infinite set. By the standard compactness argument (see Lemma 1.10), available within $\mathsf{WKL}_{0}+\Gamma$, there is a depth $d_{0}$ such that for every $f:[[0,\dots,d_{0})\cap Y]^{n}\to k$ there exists a $Z\subseteq[0,d_{0})\cap Y$ such that $\Psi(f,Z)$ and $Z$ is $(n-1)$-dense$(T(A,a),\Gamma)$.
@@
3.2 Largeness$(T)$ for Ramsey-like statements
@@
Fix $r,s\in\omega$ and a Ramsey-like-$\Pi^{1}_{2}$-statement $\Gamma\equiv(\forall f:[\mathbb{N}]^{n}\to k)(\exists Y)(Y\mbox{is infinite}\wedge\Psi(f,Y))$. A set $Z\subseteq_{fin}\mathbb{N}$ is said to be $\omega^{r}\cdot s\mbox{-large}(T,\Gamma)$ if for any $f:[Z]^{n}\to k$, there is an $(\omega^{r}\cdot s)\mbox{-large}(T)$ set $Y\subseteq Z$ such that $\Psi(f,Y)$ holds.
@@
Let $\Gamma$ be a Ramsey-like-$\Pi_{2}^{1}$-statement and $T$ be a $\Pi^{0}_{3}$ formula. Suppose $\omega^{n}$-largeness$(T,\Gamma)$ is a largeness notion provably in $\mathsf{RCA}_{0}+\mathsf{B}\Sigma^{0}_{2}+T$ for every $n\in\omega$. Then $n$-density$(T,\Gamma)$ is a largeness notion provably in $\mathsf{RCA}_{0}+\mathsf{B}\Sigma^{0}_{2}+T$ for every $n\in\omega$.
@@
Let $\Gamma$ be a Ramsey-like-$\Pi_{2}^{1}$-statement. If $\omega^{n}$-largeness$(T,\Gamma)$ is a largeness notion provably in $\mathsf{RCA}_{0}+\mathsf{B}\Sigma^{0}_{2}+T$ for every $n\in\omega$ and every $\Pi^{0}_{3}$ formula $T$, then $\mathsf{WKL}_{0}+\mathsf{B}\Sigma^{0}_{2}+\Gamma$ is a $\forall\Pi_{4}^{0}$-conservative extension of $\mathsf{RCA}_{0}+\mathsf{B}\Sigma^{0}_{2}$.
@@ 4 $\Pi^{1}_{1}$ conservation of the grouping principle
Given a model $\mathcal{M}=(M,S)\models\mathsf{RCA}_{0}$ and a set $G\subseteq M$, we write $\mathcal{M}[G]$ for the model whose first-order part is $M$ and whose second-order part consists of the $\Delta^{0}_{1}$-definable sets with parameters in $\mathcal{M}\cup\{G\}$. If $\mathcal{M}\cup\{G\}\models\mathsf{I}\Sigma^{0}_{1}$, then $\mathcal{M}[G]\models\mathsf{I}\Sigma^{0}_{1}$. By the amalgamation theorem of Yokoyama [25, Theorem 2.2] and the fact that $\mathsf{WKL}_{0}$ and $\mathsf{COH}$ are both $\Pi^{1}_{1}$ conservative over $\mathsf{RCA}_{0}+\mathsf{B}\Sigma^{0}_{2}$, Theorem 4.2 is a direct consequence of the following proposition.
@@
We are now ready to prove the main theorem of this section, namely, that the grouping principle is $\Pi^{1}_{1}$ conservative over $\mathsf{RCA}_{0}+\mathsf{B}\Sigma^{0}_{2}$. It can be alternatively proved by showing that $\Delta^{0}_{2}\mbox{-}\mathsf{GP}^{1}_{2}$ and $\mathsf{COH}$ both are $\Pi^{1}_{1}$-conservative over $\mathsf{RCA}_{0}+\mathsf{B}\Sigma^{0}_{2}$, and using the amalgamation theorem of Yokoyama [25, Theorem 2.2]. However, we give a direct argument for the sake of simplicity:
@@
The above construction shows that $\mathsf{RCA}_{0}+\mathsf{B}\Sigma^{0}_{2}$ proves the $\ll_{2}$-basis theorem for $\Delta^{0}_{2}\mbox{-}\mathsf{GP}^{1}_{2}$ (in the sense of [6]), and thus $\mathsf{GP}^{2}_{2}$ also admits $\ll_{2}$-basis theorem within $\mathsf{RCA}_{0}+\mathsf{B}\Sigma^{0}_{2}$. This implies that the above $\Pi_{1}^{1}$-conservation theorem admits poly-time proof transformation. In general, an upcoming result by Ikari and Yokoyama shows that any Ramsey-like principle provable from $\mathsf{WKL}_{0}+\mathsf{RT}^{2}_{2}$ admits poly-time proof transformation if it is $\Pi_{1}^{1}$-conservative over $\mathsf{RCA}_{0}+\mathsf{B}\Sigma^{0}_{2}$.
@@ 4.1 Finite grouping principle with $T$-apartness
5 Applications to Ramsey-like theorems

### v2

@@ Abstract
In this article, we prove that Ramsey’s theorem for pairs and two colors is a $\forall\Pi^{0}_{4}$ conservative extension of $\mathsf{RCA}_{0}+\mathsf{B}\Sigma^{0}_{2}$, where a $\forall\Pi^{0}_{4}$ formula consists of a universal quantifier over sets followed by a $\Pi^{0}_{4}$ formula. The proof is an improvement of a result by Patey and Yokoyama [22] and a step towards the resolution of the longstanding question of the first-order part of Ramsey’s theorem for pairs. For this, we introduce a new general technique for proving $\Pi^{0}_{4}$-conservation theorems.
MSC classes Primary: 03B30, 03F30 Secondary: 05D10
@@ 1 Introduction
The proof of the main theorem follows the structure of Patey and Yokoyama [22], using an indicator for $\forall\Pi^{0}_{4}$ conservation defined by Yokoyama [27, Section 4]. It is based on a variant of the notion of $\alpha$-largeness, which provides a new general technique for proving $\forall\Pi^{0}_{4}$-conservation theorems.
@@ 1.3 Structure of this paper
Then, in Section 3, we develop a framework for $\forall\Pi^{0}_{4}$ conservation of $\mathsf{RT}$-like theorems. For this, given a $\mathsf{RT}$-like theorem $\Gamma$, we first define in Section 3.1 a notion of $n$-density$(\Gamma)$, which informally asserts the existence of sufficiently large finite sets, such that $n$ consecutive applications of $\Gamma$ yield a large set. Following techniques initially defined by Kirby and Paris [14], we prove that $\mathsf{RT}^{2}_{2}$ is a $\forall\Pi^{0}_{4}$ conservative extension of $\mathsf{RCA}_{0}+\mathsf{B}\Sigma^{0}_{2}+$ “there exists $n$-dense sets for every $n\in\omega$” (see Theorem 3.9). $\forall\Pi^{0}_{4}$ conservation over $\mathsf{RCA}_{0}+\mathsf{B}\Sigma^{0}_{2}$ is then reduced to whether $\mathsf{RCA}_{0}+\mathsf{B}\Sigma^{0}_{2}$ proves the existence of these $n$-dense$(\Gamma)$ sets, for every $n\in\omega$. It is however difficult to directly prove the existence of $n$-dense$(\Gamma)$ sets, and we therefore resort in Section 3.2 to the quantitative notion of largeness, $\omega^{n}$-largeness$(T)$, to prove this existence step by step, by handling one application of $\Gamma$ at a time.
@@ 3 A framework for $\forall\Pi^{0}_{4}$ conservation
We now develop a framework for proving that $\mathsf{RT}$-like statements are $\forall\Pi^{0}_{4}$ conservative over $\mathsf{RCA}_{0}+\mathsf{B}\Sigma^{0}_{2}$. This will be divided into two sections : First, building up on the work of Kirby and Paris [14], we prove in Section 3.1 that every $\mathsf{RT}$-like theorem $\Gamma$ is $\forall\Pi^{0}_{4}$ conservative over $\mathsf{RCA}_{0}+\mathsf{B}\Sigma^{0}_{2}+$ a first-order theory stating the existence of dense$(\Gamma)$ sets. Then, in Section 3.2, we use the quantitative notion of largeness defined in Section 2 to prove the existence of dense$(\Gamma)$ sets, assuming the existence of a combinatorially simpler objects.
Given a function $f:[\mathbb{N}]^{n}\to k$ and a finite set $G=\{x_{0}<\dots<x_{\ell-1}\}$, let $f_{G}:[\ell]^{n}\to k$ be defined by $f_{G}(i_{0},\dots,i_{n-1})=f(x_{i_{0}},\dots,x_{i_{n-1}})$.
Definition 3.1 ($\mathsf{RT}$-like formula).
Given $n,k\in\omega$, an $\mathsf{RT}$-like formula is a $\Pi_{2}^{1}$-formula of the form
@@
where $\Psi(f,Y)$ is of the form $(\forall G\subseteq_{\mathtt{fin}}Y)\Psi_{0}(f_{G})$ with $\Psi_{0}$ a $\Delta^{0}_{0}$-formula.
$\mathsf{RT}$-like statements were introduced and studied in the standard realm by Patey [21], formulated in terms of forbidden patterns. The idea behind the formula $\Psi$ is that theorems from Ramsey theory are about structure and not the actual value of the domain. Thus, every solution is rewritten as a copy of the integers.
Remark 3.2.
The notion of $\mathsf{RT}$-like statement is a particular case of Ramsey-like-$\Pi_{2}^{1}$-formula, introduced by defined by Patey and Yokoyama [22]. It already encompasses Ramsey’s theorem for pairs, the Ascending Descending Sequence principle and the Erdős-Moser theorem. This section could have been formulated and proven in the more general setting of Ramsey-like-$\Pi_{2}^{1}$-formulas, with the cost of some heavier notations as in [22].
Since an $\mathsf{RT}$-like statement does not distinguish the solutions from the set of the integers, it proves a slightly stronger formula, stating that the class of solutions is dense in the partial order $([\mathbb{N}]^{\mathbb{N}},\subseteq)$.
Lemma 3.3.
Let $\Gamma\equiv(\forall f:[\mathbb{N}]^{n}\to k)(\exists Y)(Y\mbox{is infinite}\wedge\Psi(f,Y))$ be an $\mathsf{RT}$-like formula. Then $\mathsf{RCA}_{0}+\Gamma$ proves
$(\forall f:[\mathbb{N}]^{n}\to k)(\forall X\mbox{ infinite})(\exists Y\subseteq X)(Y\mbox{is infinite}\wedge\Psi(f,Y))$
Proof.
Let $f:[\mathbb{N}]^{n}\to k$ be a coloring and $X=\{x_{0}<x_{1}<\dots\}$ be an infinite set. Let $\hat{f}:[\mathbb{N}]^{n}\to k$ be defined by $\hat{f}(i_{0},\dots,i_{n-1})=f(x_{i_{0}},\dots,x_{i_{n-1}})$. By $\Gamma$, there is an infinite set $\hat{Y}\subseteq\mathbb{N}$ such that $\Psi(\hat{f},\hat{Y})$. Let $Y=\{x_{s}:s\in\hat{Y}\}$. We claim that $\Psi(f,Y)$ holds.
Say $\Psi(f,Y)\equiv(\forall G\subseteq_{\mathtt{fin}}Y)\Psi_{0}(f_{G})$ with $\Psi_{0}$ a $\Delta^{0}_{0}$-formula. Let $G\subseteq_{\mathtt{fin}}Y$, and let $\hat{G}=\{s:x_{s}\in G\}$. Say $G=\{x_{s_{0}},\dots,x_{s_{\ell-1}}\}$. Then
$f_{G}(i_{0},\dots,i_{n-1})=f(x_{s_{i_{0}}},\dots,x_{s_{i_{n-1}}})=\hat{f}(s_{i_{0}},\dots,s_{i_{n-1}})=\hat{f}_{\hat{G}}(i_{0},\dots,i_{n-1})$
hence $f_{G}=\hat{f}_{\hat{G}}$. Since $\Psi(\hat{f},\hat{Y})$ and $\hat{G}\subseteq_{\mathtt{fin}}\hat{Y}$, then $\Psi_{0}(\hat{f}_{\hat{G}})$ holds, then so does $\Psi_{0}(f_{G})$. Since it holds for every $G\subseteq_{\mathtt{fin}}Y$, then $\Psi(f,Y)$ holds. ∎
3.1 Density$(T)$ for $\mathsf{RT}$-like statements
There exists a very general way to characterize the first-order part of a second-order theory, using the notion of density, originally defined by Kirby and Paris [14]. This was later adapted by Bovykin and Weiermann [1, Theorem 1] to prove $\Pi^{0}_{2}$ conservation of combinatorial theorems, and generalized by Patey and Yokoyama [22, Theorem 3.4] for $\forall\Pi^{0}_{3}$ conservation.
@@ 3.1 Density$(T)$ for $\mathsf{RT}$-like statements
Fix an $\mathsf{RT}$-like statement
@@
Let $\Gamma$ be an $\mathsf{RT}$-like statement. Given a countable nonstandard model $\mathcal{M}\models\mathsf{I}\Sigma^{0}_{1}$ and an $M$-finite set $Z\subseteq M$ which is $c\mbox{-dense}(T,\Gamma)$ for some $c\in M\setminus\omega$, then there exists an initial segment $I$ of $M$ such that $(I,\mathsf{Cod}(M/I),a^{\mathcal{M}},A^{\mathcal{M}}\cap I)\models\mathsf{WKL}_{0}+\mathsf{B}\Sigma^{0}_{2}+\Gamma+T$ and $I\cap Z$ is infinite in $I$.
@@
Let $\Gamma$ be an $\mathsf{RT}$-like statement of the form
@@
Case $n>0$. Suppose the property to be true at rank $n-1$ and fix $Y$ an infinite set. By the standard compactness argument (see Lemma 1.10), available within $\mathsf{WKL}_{0}+\Gamma$, and by Lemma 3.3, there is a depth $d_{0}$ such that for every $f:[[0,\dots,d_{0})\cap Y]^{n}\to k$ there exists a $Z\subseteq[0,d_{0})\cap Y$ such that $\Psi(f,Z)$ and $Z$ is $(n-1)$-dense$(T(A,a),\Gamma)$.
@@
3.2 Largeness$(T)$ for $\mathsf{RT}$-like statements
@@
Fix $r,s\in\omega$ and an $\mathsf{RT}$-like-$\Pi^{1}_{2}$-statement $\Gamma\equiv(\forall f:[\mathbb{N}]^{n}\to k)(\exists Y)(Y\mbox{is infinite}\wedge\Psi(f,Y))$. A set $Z\subseteq_{fin}\mathbb{N}$ is said to be $\omega^{r}\cdot s\mbox{-large}(T,\Gamma)$ if for any $f:[Z]^{n}\to k$, there is an $(\omega^{r}\cdot s)\mbox{-large}(T)$ set $Y\subseteq Z$ such that $\Psi(f,Y)$ holds.
@@
Let $\Gamma$ be an $\mathsf{RT}$-like statement and $T$ be a $\Pi^{0}_{3}$ formula. Suppose $\omega^{n}$-largeness$(T,\Gamma)$ is a largeness notion provably in $\mathsf{RCA}_{0}+\mathsf{B}\Sigma^{0}_{2}+T$ for every $n\in\omega$. Then $n$-density$(T,\Gamma)$ is a largeness notion provably in $\mathsf{RCA}_{0}+\mathsf{B}\Sigma^{0}_{2}+T$ for every $n\in\omega$.
@@
Let $\Gamma$ be an $\mathsf{RT}$-like. If $\omega^{n}$-largeness$(T,\Gamma)$ is a largeness notion provably in $\mathsf{RCA}_{0}+\mathsf{B}\Sigma^{0}_{2}+T$ for every $n\in\omega$ and every $\Pi^{0}_{3}$ formula $T$, then $\mathsf{WKL}_{0}+\mathsf{B}\Sigma^{0}_{2}+\Gamma$ is a $\forall\Pi_{4}^{0}$-conservative extension of $\mathsf{RCA}_{0}+\mathsf{B}\Sigma^{0}_{2}$.
@@ 4 $\Pi^{1}_{1}$ conservation of the grouping principle
Given a model $\mathcal{M}=(M,S)\models\mathsf{RCA}_{0}$ and a set $G\subseteq M$, we write $\mathcal{M}[G]$ for the model whose first-order part is $M$ and whose second-order part consists of the $\Delta^{0}_{1}$-definable sets with parameters in $\mathcal{M}\cup\{G\}$. If $\mathcal{M}\cup\{G\}\models\mathsf{I}\Sigma^{0}_{1}$, then $\mathcal{M}[G]\models\mathsf{I}\Sigma^{0}_{1}$. By the amalgamation theorem of Yokoyama [26, Theorem 2.2] and the fact that $\mathsf{WKL}_{0}$ and $\mathsf{COH}$ are both $\Pi^{1}_{1}$ conservative over $\mathsf{RCA}_{0}+\mathsf{B}\Sigma^{0}_{2}$, Theorem 4.2 is a direct consequence of the following proposition.
@@
We are now ready to prove the main theorem of this section, namely, that the grouping principle is $\Pi^{1}_{1}$ conservative over $\mathsf{RCA}_{0}+\mathsf{B}\Sigma^{0}_{2}$. It can be alternatively proved by showing that $\Delta^{0}_{2}\mbox{-}\mathsf{GP}^{1}_{2}$ and $\mathsf{COH}$ both are $\Pi^{1}_{1}$-conservative over $\mathsf{RCA}_{0}+\mathsf{B}\Sigma^{0}_{2}$, and using the amalgamation theorem of Yokoyama [26, Theorem 2.2]. However, we give a direct argument for the sake of simplicity:
@@
The above construction shows that $\mathsf{RCA}_{0}+\mathsf{B}\Sigma^{0}_{2}$ proves the $\ll_{2}$-basis theorem for $\Delta^{0}_{2}\mbox{-}\mathsf{GP}^{1}_{2}$ (in the sense of [6]), and thus $\mathsf{GP}^{2}_{2}$ also admits $\ll_{2}$-basis theorem within $\mathsf{RCA}_{0}+\mathsf{B}\Sigma^{0}_{2}$. This implies that the above $\Pi_{1}^{1}$-conservation theorem admits poly-time proof transformation. In general, an upcoming result by Ikari and Yokoyama shows that any $\mathsf{RT}$-like principle provable from $\mathsf{WKL}_{0}+\mathsf{RT}^{2}_{2}$ admits poly-time proof transformation if it is $\Pi_{1}^{1}$-conservative over $\mathsf{RCA}_{0}+\mathsf{B}\Sigma^{0}_{2}$.
@@ 4.1 Finite grouping principle with $T$-apartness
5 Applications to $\mathsf{RT}$-like theorems
@@ 8 Open questions
Acknowledgement
The authors are thankful to Leszek Kołodziejczyk for insightful comments and discussions.